I need help understanding the following;
Find the LCM of the following numbers:
15,9,6

Answer by Edwin McCravy(20054)

15,9,6

Break them all down as far as they will go:

15 = 3*5
 9 = 3*3
 6 = 2*3

We look at the factor 2.  15 contains it no times. 9 contains it no times.
6 contains it 1 time.  The most number of times that any of the three
contains factor 2 is 1 time.  So the LCM will contain it 1 time. So write

LCM = 2×____

We look at the factor 3.  15 contains it 1 time. 9 contains it 2 times.
6 contains it 1 time.  The most number of times that any of the three
contains factor 2 is 2 times.  So the LCM will contain it 2 times.  So write

LCM = 2×3×3____

We look at the factor 5.  15 contains it 1 time. 9 contains it no times.
6 contains it no times.  The most number of times that any of the three
contains factor 2 is 1 time.  So the LCM will contain it 1 time.  So write

LCM = 2×3×3×5

And that's all the factors, so we multiply them all together, and get

LCD = 90

Edwin